The perimeter of a rectangular field is 660 yards. If the width is 20 yards less than the length, than what is the length?
Answer by ReadingBoosters(3246)

p = 2(l+w) = 2l+2w = 660
w= l - 20
Substitute w into the equation
660 = 2l + 2(l-20)
660 = 2l + 2l - 40
660 = 4l - 40
620 = 4l
155 yd = length